I am stumped on this question I need to evaluate this
Evaluate P(-1/2) if P(x)=2x^4+x^3+12

Answer by Alan3354(69443)

Evaluate P(-1/2) if P(x)=2x^4+x^3+12
---------
Sub -1/2 for x
P(-1/2) = 2(-1/2)^4 + (-1/2)^3 + 12
= 2*(1/16) + (-1/8) + 12
= 1/8 - 1/8 + 12
= 12